# Measuring Human Contribution in AI-Assisted Content Generation

### v1

@@ Abstract
We systematically validate the proposed method as a reliable measure of human contribution by evaluating its effectiveness, domain adaptivity, and model adaptivity. To achieve this, we construct a comprehensive dataset of AI-assisted content generation, encompassing various levels of human contribution, multiple creative domains, and outputs from different LLMs. For instance, Figure 1c illustrates the distribution of the outcomes of our proposed measure for AI-assisted poem generation, across three varying levels of human contribution, ranging from high to low, using the LLM Llama-3. Our proposed measure effectively discriminates between varying degrees of contribution, generally producing lower values for content with less human contribution. Additionally, we investigate the impact of content length, resilience to adaptive attacks, and generalization of our method in evaluation. We further apply our measure to real-world human–AI co-creation data, demonstrating its practical applicability. In brief, this paper poses a novel research question on quantitatively evaluating human contribution in AI-assisted generation and presents a simple yet effective information-based measure as a potential solution.
@@ Results
To verify the reliability of the proposed measure of human contribution, we construct a dataset of AI-assisted generation data with known varying levels of human contribution. Note that there is no absolute ground truth for assessing human contribution. Therefore, by design, our dataset spans a very large range of human contribution in AI-assisted output with distinct levels that are hardly controversial. For a comprehensive evaluation, we further vary three factors, beyond the level of human contribution: (1) domains, focusing on those where originality protection is crucial; (2) different LLMs; and (3) different random generation runs. Building this dataset primarily involves two steps: raw information collection and processing and AI-assisted content generation.
Raw Information Collection and Processing: First, we collect and process multi-level information in various domains. Specifically, we sample raw data from public datasets across the following domains: academic paper abstracts, news articles, patent abstracts, and poems. We sample 2,000 entries for each domain. For paper abstracts, each raw data entry includes content, title, and subject; for the other three domains, each raw entry includes content and title. Details of the original dataset and sampling process are provided in Supplemental Materials Section 1.1. We further process the data into a uniform structure with decreasing levels of information: content, summary, title, and subject (except poems, because of short titles), with missing parts of the raw data supplemented using GPT-3.5 [2]. The corresponding statistics are presented in Table 1.
@@
Figure 2 illustrates the human contribution results of two state-of-the-art open-source LLMs, Llama-3 and Mixtral, on the constructed dataset across various domains. From the results, we make the following observations. First, for each combination of model and data domain, varying levels of human contribution yield different distributions for measured human input in the expected direction: the lower the human author’s informational contribution in AI-assisted generation (from polishing, to generation with summary, to generation with title, and finally to generation with subject), the smaller the proposed metric’s value. For instance, when generating news articles with Llama-3, polished content demonstrates an average human contribution measurement of $85.37\%$, while content generated with a subject shows an average human contribution measurement of $30.83\%$. This indicates that our proposed measure can effectively distinguish different levels of human contribution in AI-assisted generation, providing useful measurements from an informational perspective.
@@
Second, we observe variability in the outcomes of the proposed measure across different data domains for a specific generation mode (e.g., generation with summary). For instance, the average score for content generated with a summary using Llama-3 is 55.69% for news but 84.94% for paper abstracts. These differences are reasonable because the same generation mode does not necessarily equate to a similar percentage of human contribution across different domains. As discussed in the Dataset Construction section, without an absolute ground truth, we can only generate content with approximate varying levels of human contribution. For example, the ratio of summary length to content length is significantly higher for paper abstracts than for news articles, as shown in Table 1, indicating a variation in ground truth human contribution. Therefore, in our evaluation, the primary consideration is to verify whether our measure can consistently reflect the overall pattern of varying levels of human contribution for each generation model and creative domain. This consistency would validate the reliability of our proposed measure for distinguishing different levels of human contribution.
@@
Impact of Content Length
In addition to the varying levels of human contribution present in our constructed dataset, we further validate our method by varying the length of the AI-assisted content that is generated. This helps us determine whether our method adequately evaluates the proportion of human contribution in AI-assisted generated content when the same human input information yields AI-assisted outputs of different lengths. Intuitively, when human informational input remains constant, the longer the generated content, the smaller the measured human contribution should be. To verify this, we use Llama-3 to generate AI-assisted outputs of varying lengths from titles by specifying the length of the AI-assisted output in the prompt.
Figure 3 shows the results for paper and patent abstracts with different lengths. The results for news and poems are included in Supplementary Materials Section 2.1. The results align with our initial expectation: as we require AI-assisted output to be longer, human informational contribution relative to total informational content decreases, as does our measurement of human contribution.
@@
In real-world applications, the AI model’s generative probability $p_{\bm{\theta}}$ may not be available. For instance, generative applications like ChatGPT do release generative probabilities to users. This section demonstrates whether a surrogate model with generative probability $p_{\bm{\theta}}^{\prime}$ can be employed for our assessment when the AI model’s generative probability $p_{\bm{\theta}}$ is unknown. Specifically, in this experiment, we use Llama-3 and Mistral as the surrogate models and use their generative probability $p_{\bm{\theta}}^{\prime}$ to assess the content generated by various LLMs (Llama-3, Mistral, and ChatGPT) in the constructed dataset.
Figure 5 illustrates the effectiveness of our approach in the news domain for various combinations of surrogate model (rows) and generation model (columns). Results for other domains are presented in Supplementary Materials Section 2.3. We observe that even without using the original AI model for evaluation, our proposed measure captures the expected trend in human contribution across various surrogate and generative model combinations. This validates the applicability of our measure when generation model information is unavailable. This effectiveness may be attributed to the similar generative distributions across LLMs, stemming from the universal knowledge they share during training. The gradient in human contribution across varying levels of human input is far more pronounced than the differences between the distributions themselves, indicating that our method is a robust assessment tool.
@@
Figure 6a demonstrates the overall distribution of measured human contributions across the two classes. We expect that the “editing or rewriting” class will involve more human contribution than “assisting or creative writing.” Consistent with this, the measured human contributions are generally higher for “editing or rewriting.” We present two specific cases in Figure 6b, with additional cases detailed in Supplementary Materials Section 2.4. Overall, the measured human contributions align with our expectations. For instance, the “editing or rewriting” case is measured as having 92.86% human contribution, while the “assisting or creative writing” case is measured at 16.14%. These distribution and case study results further support the validity of our method in measuring human contribution in real-world AI-assisted generation contexts.
@@ Discussion
Technological advancements in generative AI have significantly altered the content production process, resulting in the generation of a vast amount of AI-assisted content  [11, 10, 8]. This proliferation of AI-assisted generated content poses challenges to the delineation of originality of these works, sparking intense debates regarding the application of AI in areas that prioritize originality, such as education, academic research, and copyright [23, 24, 25, 31, 32, 33]. It has been increasingly recognized that defining the intellectual originality of AI-assisted generated content cannot be approached with a one-size-fits-all solution; instead, it requires consideration of the extent of human contribution to the work [20, 22, 32]. Consequently, to facilitate the generative AI era, decision makers will need credible methods to measure human contribution in AI-assisted generation across various contexts.
@@
While it serves to frame the question and provide a preliminary method for measuring human contribution, our work has several limitations. First, we currently focus on scenarios where human input is available and reliable for evaluation. Further research is necessary to measure human contributions in scenarios where human input is unknown. Second, the current framework focuses on textual output from LLMs. However, originality issues related to AI-assisted generation are not limited to text; they also extend to images, audio, video, computer code, etc. Incorporating non-textual output raises even more complex problems due to the change in modality between human input and AI-assisted output. We aim to explore originality issues in such scenarios in future research. Third, this research does not include experiments which entailing multiple rounds of human-AI interaction, prompting and content generation. In the more complex situation where content is generated with multiple prompting rounds, human contribution should be measured by considering the human’s intent and understanding of the generated content, which may inform the contribution assessment. Fourth, caution is warranted when measuring human contribution in AI-assisted content generation in the copyright domain. Human edition, selection, and compilation of AI-assisted content may provide significant creative input, which could be relevant when assessing authorship for copyright purposes.
@@
The objective of this study is to pose a research question and propose a framework for measuring human contribution in AI-assisted content generation. This question and framework aim to facilitate originality delineation in the era of creation with the assistance of AI. Simultaneously, this work seeks to inspire more research on technical methods that can support the enhancement of relevant regulations in the context of widespread AI utilization in various scenarios. A potential risk is that in real-world applications of the proposed framework, there might be targeted adaptive attacks aimed at manipulating the results to artificially elevate the assessed level of human contribution. Although this paper examines two adaptive attacks and verifies the robustness of the proposed measure against them, more sophisticated and advanced attacks may arise in real-world scenarios. We hope to further understand and mitigate such risks in future work.
The authors of this paper introduce a mere method to technically measure the human contribution in AI-assist content generation which can be potentially used in various scenarios. However, the paper does not intend to discuss the complex copyright legal and policy issues related to “originality” or “eligibility,” nor it reflect any of Microsoft’s legal and policy positions on the copyright issues.
@@ Methods
In this section, we first introduce related concepts derived from information theory [27]; we then provide our definition of human contribution. In the following definition, we consider an AI model $M_{\theta}$, its generative distribution $p_{\bm{\theta}}$, human input $\bm{x}$, and AI-assisted output $\bm{y}$. First, we quantify the information content within the generated output $\bm{y}$ through the concept of self-information. Self-information measures the level of surprisal associated with the outcome of a random variable, which is related to the probability of that outcome occurring. In this context, content that is less probable in its generation is deemed more informative. We represent the self-information of the generated output $\bm{y}$ as follows:
$I(\bm{y})=-log(p_{\bm{\theta}}(\bm{y})),$ (1)
where $p_{\bm{\theta}}(\bm{y})$ is the probability that the content $\bm{y}$ is generated without any condition.
On the other hand, when conditioned on human input $\bm{x}$, the information content within the generated output $\bm{y}$ transforms into conditional self-information. Conditional self-information quantifies the information contained in an outcome of a random variable, given the occurrence of another event. Here, we represent the conditional self-information of the generated output $\bm{y}$ given the human input $\bm{x}$ as follows:
$I(\bm{y}|\bm{x})=-log(p_{\bm{\theta}}(\bm{y}|\bm{x})),$ (2)
where $p_{\bm{\theta}}(\bm{y})$ is the probability that the content $\bm{y}$ is generated conditioned on human input $\bm{x}$.
Based on these two information concepts, we define the mutual information between the generated content $\bm{y}$ and the human input $\bm{x}$ as the information gain during generation when the human input is known. This signifies the reduction in surprisal when human input $\bm{x}$ is for generating content $\bm{y}$ is provided, defined as follows:
$I(\bm{x};\bm{y})=I(\bm{y})-I(\bm{y}|\bm{x}).$ (3)

### v2

@@ Abstract
We systematically validate the proposed method as a reliable measure of human contribution by evaluating its effectiveness, domain adaptivity, and model adaptivity. To achieve this, we construct a comprehensive dataset of AI-assisted content generation, encompassing various levels of human contribution, multiple creative domains, and outputs from different LLMs. For instance, Figure 1c illustrates the distribution of the outcomes of our proposed measure for AI-assisted poem generation, across three varying levels of human contribution, ranging from high to low, using the LLM Llama-3. Our proposed measure effectively discriminates between varying degrees of contribution, generally producing lower values for content with less human contribution. Additionally, we investigate the impact of content length, resilience to adaptive attacks, and generalization of our method in evaluation. We further apply our measure to real-world human–AI co-creation data, demonstrating its practical applicability. To broaden the applicability of our human contribution evaluation framework to real-world scenarios—where humans may engage in complex content generation processes or where human input is unknown or impossible to measure—we propose an extension that estimates human contribution based on any text alone, without requiring access to its generation process. In brief, this paper poses a novel research question on quantitatively evaluating human contribution in AI-assisted generation and presents simple yet effective information-based measures as potential solutions.
@@ Results
To verify the validity and reliability of the proposed measure of human contribution, we construct a dataset of AI-assisted generation data with known varying levels of human contribution. Note that obtaining a dataset with clear, real-valued labels poses a significant challenge, as there is no established method for quantifying human contribution when both human input and output are known. This gap motivates us to propose our conceptual framework as a meaningful first step toward addressing this complex issue. Therefore, by design, our dataset spans a very large range of human contribution in AI-assisted output with distinct levels that are hardly controversial. For a comprehensive evaluation, we further vary three factors, beyond the level of human contribution: (1) domains, focusing on those where originality protection is crucial; (2) different LLMs; and (3) different random generation runs. Building this dataset primarily involves two steps: raw information collection and processing and AI-assisted content generation.
Raw Information Collection and Processing: First, we collect and process multi-level information in various domains. Specifically, we sample raw data from public datasets across the following domains: academic paper abstracts, news articles, patent abstracts, and poems. We sample 2,000 entries for each domain. For paper abstracts, each raw data entry includes content, title, and subject; for the other three domains, each raw entry includes content and title. Details of the original dataset and sampling process are provided in Supplemental Materials Section 1.1. We further process the data into a uniform structure with decreasing levels of information: content, summary, title, and subject (except poems, because of less-informative titles), with missing parts of the raw data supplemented using GPT-3.5 [2]. The corresponding statistics are presented in Table 1.
@@
Overall Trend Analysis: Figure 2 illustrates the human contribution results of two state-of-the-art open-source LLMs, Llama-3 and Mixtral, on the constructed dataset across various domains. From the results, we make the following observations. First, for each combination of model and data domain, varying levels of human contribution yield different distributions for measured human input in the expected direction: the lower the human author’s informational contribution in AI-assisted generation (from polishing, to generation with summary, to generation with title, and finally to generation with subject), the smaller the proposed metric’s value. This indicates that our proposed measure can effectively distinguish different levels of human contribution in AI-assisted generation, providing useful measurements from an informational perspective.
@@
Second, we observe variability in the outcomes of the proposed measure across different data domains for a specific generation mode (e.g., generation with summary). These differences are reasonable because the same generation mode does not necessarily equate to a similar percentage of human contribution across different domains. For example, in our dataset, the News category consists of full articles with detailed and vivid descriptions, while Paper Abstracts are typically concise paragraph. However, the summaries, titles, and subjects used for generation in our dataset for both categories are of comparable information content. Consequently, when generating a news article versus a paper abstract from a summary, title, or subject, the expected human contribution distributions differ due to the inherent difference in the information content of the data. Therefore, in our evaluation, the primary consideration is to verify whether our measure can consistently reflect the overall pattern of varying levels of human contribution for each generation model and creative domain. This consistency would validate the reliability of our proposed measure for distinguishing different levels of human contribution.
Human-annotation Dataset: To further validate the reliability and consistency of our quantitative analysis, we construct a dataset with human annotations. As discussed earlier, obtaining a dataset with clear, real-valued labels remains a significant challenge, as there is no established method for quantifying human contribution when both human input and output are known. Inspired by evaluation platforms such as Chatbot Arena [29], we collect human annotation data by pairwise comparisons. Specifically, we sample 1,000 pairs of input-output data points generated by Llama-3 across all four domains and categories with the condition that the measured human contribution gap between the two points exceeds 0.1 using our method. This threshold is chosen to ensure the meaningfulness of the comparisons and to reduce ambiguity during annotation. Pairs of input-output data points are so annotated as to reflect which one contains a higher degree of human contribution (annotation instructions are detailed in Supplementary Materials section 1.3). Using this annotated dataset, we evaluate how well the relative magnitudes derived from our measurements align with human judgments in these comparisons. The results indicate that 97.4% of the data exhibit consistency between our measurements and human annotations. Moreover, as shown in Supplementary Figure 2, the inconsistent pairs are predominantly concentrated in cases where the measured human contribution gap between two data points is less than 0.2. This further lends support to the robustness and validity of our measurement, demonstrating its consistency across different domains and setups.
Controlled Experiments: In addition to the varying levels of human contribution present in our constructed dataset, we further validate our method by manipulating the total information content of the output while controlling the information content of the input. By doing so, we aim to verify whether our measurement exhibits the expected property that, when the input is controlled, increasing the output information leads to a lower human contribution (as defined by the ratio of human-contributed information relative to the total information in the output). To achieve this, we adjust the output length requirements in the input prompts while keeping the other input information constant. We assume that much longer outputs generally have higher information content. Specifically, we use Llama-3 to generate AI-assisted outputs of varying lengths from titles by specifying the desired output length in the prompt. Figure 3 shows the results for paper and patent abstracts with different lengths, while the results for news and poems are included in Supplementary Materials Section 2.4. The results align with our expectation: as the AI-assisted output contains more information content, the proportion of human informational contribution relative to the total informational content decreases, as reflected by our measurement of human contribution.
@@
Figure 4: The distribution of the proposed measurement outcomes for academic paper and patent abstracts, generated using different temperature settings with Llama-3. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the results align with our expectation that larger temperature leads to smaller measured human contribution.
Impact of Generative Model Temperature
Temperature refers to the parameter that controls the randomness of output in a generative model. During the generation process, a lower temperature makes the model more deterministic by favoring high-probability tokens, while a higher temperature introduces greater variability, resulting in more diverse but less predictable outputs. In this section, we study the impact of generative model temperature on the outcomes of our measurement. While the “human effort” (human input) remains unchanged across different temperature settings, we expect human contribution—defined as the proportion of informational content in the output attributable to human input—to decrease as the temperature increases. This is because the output becomes less dependent on the input information and increasingly reflects the model’s explorative generation. In extreme cases, when the temperature approaches infinity, the output of the language model would become entirely independent of the human input, resulting in an expected human contribution of zero.
Figure 4 demonstrates the comparison between temperature settings from 0.3 to 0.9 during the generation with Llama-3 for paper and patent abstracts, while the results for news and poems are included in Supplementary Materials Section 2.5. We can observe that generally, a higher temperature leads to a smaller measured human contribution across different generation categories. This validates that our measurement accurately reflects the diminishing influence of human input to AI-assisted output with increasing temperature of the generative model.
@@
In real-world applications, the AI model’s generative probability $p_{\bm{\theta}}$ may not be available. For instance, generative applications like ChatGPT do release generative probabilities to users. This section demonstrates whether a surrogate model with generative probability $p_{\bm{\theta}}^{\prime}$ can be employed for our assessment when the AI model’s generative probability $p_{\bm{\theta}}$ is unknown. Specifically, in this experiment, we use Llama-3 and Mistral as the surrogate models and use their generative probability $p_{\bm{\theta}}^{\prime}$ to assess the content generated by various LLMs (Llama-3, Mistral, and ChatGPT) in the constructed dataset. Note that generation models can include chatbots or models that do not release their generative probabilities, while surrogate models require access to generative probability distributions for evaluation.
Figure 6 illustrates the effectiveness of our approach in the news domain for various combinations of surrogate model (rows) and generation model (columns). Results for other domains are presented in Supplementary Materials Section 2.7. We observe that even without using the original AI model for evaluation, our proposed measure captures the expected trend in human contribution across various surrogate and generative model combinations. This validates the applicability of our measure when generation model information is unavailable. This effectiveness may be attributed to the similarity in generative distributions across LLMs, stemming from the universal knowledge they share during training. The gradient in human contribution across varying levels of human input is far more pronounced than the differences between the distributions themselves, indicating that our method is a robust assessment tool.
@@
Figure 7a demonstrates the overall distribution of measured human contributions across the two classes. We expect that the “editing or rewriting” class will involve more human contribution than “assisting or creative writing.” Consistent with this, the measured human contributions are generally higher for “editing or rewriting.” We present two specific cases for two classes respectively in Figure 7b, with additional cases detailed in Supplementary Materials Section 2.8. Overall, the measured human contributions align with our expectations. For instance, the “editing or rewriting” case is measured as having 92.86% human contribution, while the “assisting or creative writing” case is measured at 16.14%. These distribution and case study results further support the validity of our method in measuring human contribution in real-world AI-assisted generation contexts.
Figure 8: The distribution of the outcomes of the proposed estimation for the constructed dataset. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the proposed estimation exhibits the expected trend that lower values are obtained for the generated content with less human contribution.
Human Contribution Estimation Without Human Input
To better address real-world scenarios where the AI-assisted generation process is unknown, we extend our method to estimate the human contribution of any given text alone without knowledge of human input and output generation processes. Clearly, we have less information now than the situation considered before. This loss of information leads to a degree of uncertainty, as the same output can be generated from inputs with varying levels of contribution. However, we address this problem of uncertainty by redefining our goal to estimate the proportion of information content in the AI-assisted output that is necessarily provided by humans. This is also referred to as the minimum human contribution for human input $\bm{x}$ to plausibly generate the AI-assisted output $\bm{y}$. This focus on minimal human contribution stems from several practical considerations, such as evaluating the unique original informativeness value that LLMs cannot easily generate, as detailed in the Methods section.
To compute this, we adapt the original formula $\frac{I(\bm{x};\bm{y})}{I(\bm{y})}$. While the denominator $I(\bm{y})$ can still be directly calculated, the numerator $I(\bm{x};\bm{y})=I(\bm{y})-I(\bm{y}|\bm{x})$ is estimated as its minimum possible value for a valid hypothetical $\bm{x}$ that plausibly generates $\bm{y}$. Intuitively, if a hypothetical $\bm{x}$ plausibly generates $\bm{y}$, the likelihood of $\bm{y}$ being generated based on $\bm{x}$, $p_{\theta}(\bm{y}\mid\bm{x})$, should be sufficiently high (have a lower bound). This implies that the conditional information (or surprisal) $I(\bm{y}|\bm{x})=-log(p_{\theta}(\bm{y}\mid\bm{x}))$ of generating $\bm{y}$ given $\bm{x}$ should have an upper bound. We leverage this upper bound to compute the minimum possible value of $I(\bm{x};\bm{y})$. The detailed motivation for estimating minimum human contribution, along with the calculation method, is provided in the Methods section.
Overall Trend Analysis: In Figure 8, we present the human contribution estimation results for two state-of-the-art open-source LLMs, Llama-3 and Mixtral, applied to the constructed dataset across various domains. We make the following observations: First, the proposed estimation method produces results that follow the expected trend in our dataset: as the human author’s informational contribution decreases during AI-assisted generation, the estimated metric value correspondingly decreases. Second, as we suppress the diverse real values of $I(\bm{y}\mid\bm{x})$ to an estimated boundary value that is empirically chosen, the resulting distributions do not fully match human contribution evaluations where $\bm{x}$ is known (Figure 2). Finally, since this is an estimation process, there are cases where values fall below zero. These can be interpreted as instances where the corresponding output might plausibly be generated with minimal or negligible human contribution under the minimal human contribution assumption. Overall, our method provides a meaningful solution for estimating human contribution in situations where direct input information and generation processes are unavailable.
Figure 9: a. Illustration of the tested multi-round refinement scenario. The human contribution to the output is expected to decrease from round 1 to round 3, as the output progressively incorporates more information with minimal additional input information from the human. b. The evaluated human contribution relationship across outputs from different rounds. Our evaluation produces results that align with our expectations: the human contributions evaluated from the outputs of the later rounds are consistent with those of the corresponding previous rounds, showing a gradual decrease from round to round.
Multi-Round Generation Analysis: In addition, we further investigate the validity of our evaluation in the context of multi-round refinement. We design an experiment, as illustrated in Figure 9a, where we sample the generation of 1,000 news articles from the original dataset for the first round of generation (covering all generation modes). Subsequently, we add second and third rounds of refinement, requiring the model to continuously incorporate additional information. Intuitively, the proportion of human contribution decreases from the round 1 output to the round 3 output, as the output progressively incorporates more information while minimal additional input from the human is given.
The experimental results are presented in Figure 9b. The left panel shows the relationship between the human contribution evaluated for the round 1 output (x-axis) and the round 2 output (y-axis), while the right panel depicts the relationship between the round 2 output (x-axis) and the round 3 output (y-axis). These figures illustrate changes in human contribution across different rounds of output. The results align with our expectations: there is consistency across rounds, meaning the human contribution differences introduced by various generation modes (e.g., polish, generation with summary, title, or subject) in the first round persist in the subsequent rounds. At the same time, most data points lie below the line $y=x$, indicating that human contribution generally decreases in the later rounds. Statistically, 87.12% of cases show a reduction in human contribution from round 1 to round 2 , and 94.37% show a reduction from round 2 to round 3.
@@ Discussion
Technological advancements in generative AI have significantly altered the content production process, resulting in the generation of a vast amount of AI-assisted content  [11, 10, 8]. This proliferation of AI-assisted generated content poses challenges to the delineation of originality of these works, sparking intense debates regarding the application of AI in areas that prioritize originality, such as education, academic research, and copyright [23, 24, 25, 32, 33, 34]. It has been increasingly recognized that defining the intellectual originality of AI-assisted generated content cannot be approached with a one-size-fits-all solution; instead, it requires consideration of the extent of human contribution to the work [20, 22, 33]. Consequently, to facilitate the further development of generative AI, decision makers will need credible methods to measure human contribution in AI-assisted generation across various contexts.
@@
While it serves to frame the question and provide a preliminary method for measuring human contribution, our work has several limitations. First, the current framework focuses on textual output from LLMs. However, originality issues related to AI-assisted generation are not limited to text; they also extend to images, audio, video, computer code, etc. Incorporating non-textual output raises even more complex problems due to the change in modality between human input and AI-assisted output. We aim to explore originality issues in such scenarios in future research. In other modalities, since data may no longer be discrete, this framework may not be directly applicable. However, as generative models are fundamentally grounded in certain generative probabilistic distributions, we hope that the information theory-based framework proposed in this work can serve as a foundation and inspiration for addressing originality challenges across diverse modalities. Second, caution is warranted when measuring human contribution in AI-assisted content generation in the copyright domain. Human edition, selection, and compilation of AI-assisted content may provide significant creative input, which could be relevant when assessing authorship for copyright purposes. Finally, we acknowledge that “human contribution” can have different meanings. In this work, we focus specifically on measuring informational contribution, while leaving other interpretations, such as prompt-engineering effectiveness, for future exploration.
@@
The objective of this study is to pose a research question and propose a framework for measuring human contribution in AI-assisted content generation. This question and framework aim to facilitate originality delineation in the era of creation with the assistance of AI. Simultaneously, this work seeks to inspire more research on technical methods that can support the enhancement of relevant regulations in the context of widespread AI utilization in various contexts. A potential risk is that in real-world applications of the proposed framework, there might be targeted adaptive attacks aimed at manipulating the results to artificially elevate the assessed level of human contribution. Although this paper examines two adaptive attacks and verifies the robustness of the proposed measure against them, more sophisticated and advanced attacks may arise in real-world scenarios. We hope to further understand and mitigate such risks in future work.
The authors of this paper introduce a method to technically measure human contribution in AI-assist content generation. The method is widely applicable in different situations. However, the paper does not intend to discuss the complex copyright legal and policy issues related to “originality” or “eligibility,” nor it reflect any of Microsoft’s legal and policy positions on the copyright issues.
@@ Methods
In this section, we first introduce related concepts derived from information theory [27]; we then provide our definition of human contribution. In the following definition, we consider an AI model $M_{\theta}$, its generative distribution $p_{\theta}$, human input $\bm{x}$, and AI-assisted output $\bm{y}$. First, we quantify the information content within the generated output $\bm{y}$ through the concept of self-information. Self-information measures the level of surprisal associated with an event, reflecting the probability of that event occurring. In this context, generating content that is less probable is considered more informative event. We represent the self-information of the generated output $\bm{y}$ as follows:
$I(\bm{y})=-log(p_{\theta}(\bm{y})),$ (1)
where $p_{\theta}(\bm{y})$ is the probability that the content $\bm{y}$ is generated without any condition.
On the other hand, when conditioned on human input $\bm{x}$, the information content within the generated output $\bm{y}$ transforms into conditional self-information. Conditional self-information quantifies the information contained in an event, given the occurrence of another event. Here, we represent the conditional self-information of the generated output $\bm{y}$ given “human input is $\bm{x}$” as follows:
$I(\bm{y}\mid\bm{x})=-log(p_{\theta}(\bm{y}\mid\bm{x})),$ (2)
where $p_{\theta}(\bm{y}\mid\bm{x})$ is the probability that the content $\bm{y}$ is generated conditioned on human input $\bm{x}$. It is noteworthy that both information and conditional information are calculated for the specific event “generated output is $\bm{y}$,” rather than the entire generative distribution—which would be difficult to measure due to the vast space of all possible generated contents. This allows the probabilities $p_{\theta}(\bm{y})$ and $p_{\theta}(\bm{y}\mid\bm{x})$ to be easily computed using the corresponding generative probabilities, without and with human input. Specifically, considering the generative process of the autoregressive LLM $M_{\theta}$, these probabilities can be computed by multiplying the probabilities of each token $y_{i}$ being sampled from the generative distribution, conditioned on the previously generated tokens (and the human input $\bm{x}$, if provided):
$p_{\theta}(\bm{y})=\prod_{i=1}^{N}p_{\theta}(y_{i}\mid\bm{y}_{<i}),$ (3)
$p_{\theta}(\bm{y}\mid\bm{x})=\prod_{i=1}^{N}p_{\theta}(y_{i}\mid\bm{y}_{<i},\bm{x}),$ (4)
where $N$ is the total number of tokens in $\bm{y}$..
Based on these two information concepts, we define (point-wise) mutual information between the generated content $\bm{y}$ and the human input $\bm{x}$ as the reduction in surprisal when human input $\bm{x}$ is known for generating content $\bm{y}$. This indicates how much information content in $\bm{y}$ can be attributed to human input $\bm{x}$. Formally, it is defined as:
$I(\bm{x};\bm{y})=I(\bm{y})-I(\bm{y}\mid\bm{x}).$ (5)